# A Survey on Large Language Models with Multilingualism: Recent Advances and New Frontiers

### v1

@@ 6.2.1 Legal LLMs in Multilingual Scenarios
Similar to the medical domain, the applications of LLMs in the legal domain principally center on English. Several precedent attempts such as Chatlaw Cui et al. (2023b), Lawyer LLaMA Huang et al. (2023f), SaulLM Colombo et al. (2024), and LegalBERT Chalkidis et al. (2020) expand the general ability of LLMs to the legal domain. The universal performance degradation has been observed when expanding to other languages.
To address the specific problems in the legal domain, the proposed models need to adapt the legal features, which are factual, ambiguous, structured, and timely Bansal et al. (2019); Gordon et al. (2009), compared to other domains. Brugger et al. Brugger et al. (2023) take a preliminary step in the multilingual legal Sentence Boundary Detection (SBD) covering 6 languages. Christen et al. Christen et al. (2023) conduct similar research on the multilingual Negative Scope Resolution (detecting words affected by negation cue) task and Baumgartner et al. Baumgartner et al. (2024) extend legal judgment prediction to German, French and Italian. Furthermore, for a comprehensive evaluation, Niklaus et al. Niklaus et al. (2023a) collect 11 natural language understanding legal datasets covering a total of 24 languages and 8 subdivisions, which is the first cross-lingual legal benchmark (LEXTREME) showing there is still improvement room even for the popular models like ChatGPT. After the ChatGPT series models were proposed, Nguyen et al. Nguyen et al. (2024) implement a preliminary empirical comparison between ChatGPT and GPT-4 on the “COLIEE” benchmark Rabelo et al. (2020, 2022) encompassing Japanese and English, where GPT-4 consistently surpasses its predecessor but still falls behind human performance.
For a more in-depth exploration of LLMs in multilingual legal issues, Niklaus et al. Niklaus et al. (2023b) construct a multilingual legal-domain corpus with 689GB, MultiLegalPile, whose detail is shown in Table 9. They train two PLMs based on XLM-R Conneau et al. (2020b) and the Longformer Beltagy et al. (2020) with this corpus, achieving the state-of-the-art performance on the LEXTREME Niklaus et al. (2023a) compared to XLM-R in most languages. Moreover, the English version of the model achieves state-of-the-art performance on 5 out of 7 tasks in LexGLUE Chalkidis et al. (2022b), underscoring the exceptional cross-lingual ability and legal-domain expertise. Trautmann et al. Trautmann et al. (2022) focus on employing legal prompt engineering (LPE) to enhance the capabilities of LLMs, thereby mitigating the challenges posed by the scarcity of cross-lingual legal data and the substantial computational resources required. Although improvements over baselines are observed, bridging the performance gap to match the supervised methods remains a significant endeavor.
@@ 6.3 Limitations and Future Directions on Multi-Domain
Data scarcity and translation issues. A powerful multilingual LLM, especially in specific domains, is predominantly hindered by the scarcity of domain data. Although knowledge transfer provides some relief, the issue of under-representation persists, particularly for low-resource languages, and becomes further compounded when extending to specific domains. Machine translation techniques offer a potential solution to mitigate this issue, however, they lead to new challenges Nicholas and Bhatia (2023). On one hand, machine translation systems introduce errors, particularly when handling domain-specific terminology across multiple languages. Terms or phrases that native speakers do not use may also be included in the translated corpus. On the other hand, the models suffer from comprehensively understanding and accounting for the local and cultural context of the target language, complicating the task of in-depth and high-level feature learning and capturing.
@@ 7 Multilingual Data Resource
Amazon intent mteb (a) 2022 49 2.02 GB Amazon Amazon
Amazon reviews Keung et al. (2020a) 2020 6 - Amazon Amazon
Bactrian-x Li et al. (2023d) 2023 51 - - MBZUAI
Biblenlp eBible.org 2024 861 581 MB Bible -
Bloom-lm Le Scao et al. (2022) 2022 364 - - SIL International
CC100 Conneau et al. (2020a) Wenzek et al. (2020) 2020 109 185 GB CC Facebook
CulturaX Nguyen et al. (2023) 2023 167 27 TB mC4, OSCAR University of Oregon
GPT-4 Prompts erfanzar 2024 5 1.05 GB GPT-4 -
Guanaco Dettmers et al. (2023) 2023 4 400 MB - University of Washington
HPLT de Gibert et al. (2024) 2023 75 22 TB CC, Internet Archive HPLT project
IWSLT 2017 Cettolo et al. (2017) 2017 10 4.24 GB TED Talks FBK
mC4 Raffel et al. (2019) 2019 101 9.7 TB CC Google
Mewsli-x Ruder et al. (2021a) 2021 50 285 MB WikiNews and Wikipedia DeepMind
Minds14 Gerz et al. (2021) 2021 12 471 MB Banking Assistant PolyAI Limited
Miracl Zhang et al. (2022b) 2023 18 - Wikipedia University of Waterloo
MLDR Chen et al. (2024b) 2024 13 - Wikipeida, Wudao, mC4 BAAI
MMedC Qiu et al. (2024b) 2024 6 105 GB Multiple SJTU
MQA De Bruyn et al. (2021) 2021 36 122 GB CC CLiPS Research
Multi-sentiments tyqiangz 2022 12 141 MB Multiple Sources -
Multiconer Fetahu et al. (2023a) Fetahu et al. (2023b) 2023 12 338 MB Wikipedia Amazon
Nomiracl Thakur et al. (2023) Thakur et al. (2024) 2023 16 - Wikipedia University of Waterloo
Open Subtitles Lison and Tiedemann (2016) 2023 58 273 MB Subtitles -
OSCAR Abadji et al. (2022) 2020 166 6.3 TB CC University of Orego
Para-pat Soares et al. (2020) 2020 15 2.57 GB Google Cloud University of Sheffield
Project Gutenberg Gutenberg 2023 11 14.4 GB eBook Project Gutenberg
ShareGPT RyokoAI 2023 5 1.08 GB GPT / Human Ronso
SREDFM Huguet Cabot et al. (2023) 2023 18 8.29 GB Wikidata, Wikipedia Babelscape
TED Talks Qi et al. (2018) 2018 55 - TED Talks CMU
TED-talks-iwslt Cettolo et al. (2012) 2012 104 25 GB TED Talks FBK
Toxi-text FredZhang 2023 55 1.96 GB - -
UD Nivre et al. (2020) 2023 102 2.19 GB - Universal Dependencies
Wikiann Pan et al. (2017) Rahimi et al. (2019) 2019 173 143 MB Wikipedia RPI
Wikipedia Foundation 2024 322 71.8 GB Wikipedia Wikimedia Foundation
Wit Base Srinivasan et al. (2021) 2021 105 5.15 GB Wikipedia Google
xP3 Muennighoff et al. (2022a) 2022 277 - - Cohere For AI
@@
LLMs are data-driven, thus the impressive learning capabilities of LLMs stem from their massive model sizes and extensive training datasets, which have been proven in high-resource languages. However, English stands as the closest approximation to the lingua franca, wielding dominance across various domains. With the largest number of total speakers, its prominence extends far and wide Pereltsvaig (2020), where English reigns as the primary language of the internet Schneider (2018); Pennycook (2002). Meanwhile, English is the main language used in the higher economic status countries of the world, such as America, Britain, and other Western countries Joshi et al. (2020). Therefore, existing data resources focus on the English-centric, which came at the expense of regional and indigenous languages, exacerbating language endangerment and economic marginalization Dodge et al. (2021). Due to the lack of resources, this situation deeply restricts the development of multilingual models, and it is a vicious circle Caswell et al. (2020). Moreover, low-resource languages suffer from lower quality, due to mislabeling or inadequate representation of native usage. This situation is especially prevalent with web-crawled data which predominantly consists of pornographic, nonsensical, or non-linguistic content Kreutzer et al. (2022).
@@
As shown in Table 10, we collect as much large-scale multilingual data resource as is reliable. We can observe that the scale of existing multilingual resources is much smaller than that of English monolingual resources (only four datasets with the TB level). And the low-/medium-resource languages typically derive data from a narrower range of sources compared to their high-resource counterparts Nekoto et al. (2020). The available data mainly originates from sources, such as Wikipedia, the Bible, and Common Crawl. In addition, these data suffer from bias and fairness, which we will present in Section 9. For instance, a gender bias exists in Wikipedia, with studies revealing a persistently low percentage of women editing articles Callahan and Herring (2011). The vicious circle of multilingual data includes the issue of open source issues. Due to the high cost of high-quality multilingual resources, researchers and companies are reluctant to share resources in the open-resource community (i.e., “data island”) Peris et al. (2023); Yao et al. (2024); Yan et al. (2024); Kandpal et al. (2022). This situation results in multilingual research staying at the data level and ignoring the competition in the model paradigm. The united governments, companies, and researchers must start a virtuous cycle of multilingual data resources. Access to abundant, meticulously collected datasets in a language empowers researchers and developers to construct models and benchmarks. The abundance of models and benchmarks, in turn, fosters increased publication, facilitates communication, and promotes real-world application scenarios for companies. These outputs have the potential to attract more users, while government-mandated guidelines help generate non-toxic data, which can be repurposed for further research and development.
@@ 8 Multilingual Benchmark and Evaluation
Afrisent Muhammad et al. (2023) 2023 14 $\times$ Sentiment Analysis U.Porto
ASPEN Razumovskaia et al. (2024) 2022 31 Story Planning Cambridge University
BELEBELE Bandarkar et al. (2023) 2023 122 QA Meta
BioInstructQA BioMistral 2024 7 Question Answering Avignon Université
Bucc-bitext-mining mteb (b) 2022 5 $\times$ Mixed HuggingFace
Cross-Sum Bhattacharjee et al. (2023) 2021 45 Summarization BUET
Crossmodal-3600 Thapliyal et al. (2022) 2022 36 Image Captioning Google
Exams Hardalov et al. (2020) 2020 15 Examination Sofia University
Fairlex Chalkidis et al. (2022a) 2022 5 $\times$ Fairness University of Copenhagen
FLORES-200 Team (2022) Goyal et al. (2022) Guzmán et al. (2019) 2022 200 Machine Translation Meta
GEOMLAMA Yin et al. (2022) 2022 5 Knowledge Diversity University of California
Humaneval-XL Peng et al. (2024) 2024 23 Code Generation University of Copenhagen
M-Hellaswag Dac Lai et al. (2023) 2023 35 Commonsense NLI University of Oregon
M-MMLU Dac Lai et al. (2023) 2023 35 NLU University of Oregon
M3Exam Zhang et al. (2023d) 2023 9 $\times$ Examination DAMO Academy
M3LS Verma et al. (2023) 2023 20 $\times$ Summarization IITs
MARC Keung et al. (2020b) 2020 6 $\times$ Text Classification Amazon
MasakhaNER Adelani et al. (2021) 2021 10 Named Entity Recognition Masakhane
Masakhanews Adelani et al. (2023a) 2023 16 $\times$ News Topic Classification Masakhane
MASSIVE FitzGerald et al. (2022) 2022 51 Intent Recognition, Slot Filling Amazon
MaXM Changpinyo et al. (2023) 2023 7 Visual Question Answering Google
MEGA Ahuja et al. (2023) 2023 70 $\times$ Mixed UW
MEGAVerse Ahuja et al. (2024) 2023 83 $\times$ Mixed Microsoft
Mela Zhang et al. (2024b) 2022 10 $\times$ Linguistic Acceptability SJTU
MGSM Shi et al. (2022a) 2022 11 Mathematical Reasoning Google
MLQA Lewis et al. (2020b) 2020 7 Question Answering Facebook
MMedBench Qiu et al. (2024c) 2023 6 $\times$ Medical SJTU
Multi-CoNER Malmasi et al. (2022) 2022 11 $\times$ Named Entity Recognition Amazon
MULTIEURLEXDOC Chalkidis et al. (2021) 2021 23 Legal Topic Classification University of Copenhagen
Multilingual-Fig-QA Kabra et al. (2023) 2023 7 $\times$ NLI CMU
NusaX Winata et al. (2023b) 2023 10 Mixed Bloomberg
ODEX Wang et al. (2023d) 2022 14 Code Generation CMU
OPUS-100 Zhang et al. (2020) 2020 100 Machine Translation University of Edinburgh
Paws-X Yang et al. (2019) 2019 7 Paraphrase Identification Google
Pmindiasum Urlana et al. (2023) 2023 14 Summarization IIIT Hyderabad
PRESTO Goel et al. (2023) 2023 6 $\times$ Conversational Parsing University of Rochester
SEAHORSE Clark et al. (2023) 2023 6 $\times$ Summarization DeepMind
Sib200 Adelani et al. (2023b) 2024 205 Topic Classification UCL
SMiLER Seganti et al. (2021) 2021 14 $\times$ Entity and Relation Extraction Samsung
STSB-multi-mt May (2021) 2021 10 Text Similarity -
Tatoeba-mt Tiedemann (2020) 2020 93 Machine Translation Tatoeba.org
Tydip Srinivasan and Choi (2022) 2022 9 $\times$ Identify Politeness Levels UT-Austin
TyDiQA Clark et al. (2020) 2020 11 $\times$ Question Answering Google
Universal Dependencies Nivre et al. (2020) 2022 146 $\times$ Parser -
X-CLAIM Mittal et al. (2023) 2023 6 $\times$ Realworld Claims MBZUAI
X-RiSAWOZ Moradshahi et al. (2023) 2023 6 Dialogue Utterances Stanford University
XCOPA Ponti et al. (2020) 2020 12 Commonsense Reasoning Cambridge University
XCSQA Lin et al. (2021b) 2021 11 Commonsense Reasoning USC
xDial-Eval Zhang et al. (2023e) 2023 10 Dialogues NUS
XGLUE Liang et al. (2020) 2020 - $\times$ Mixed Microsoft
XL-SUM Hasan et al. (2021) 2021 44 $\times$ Summarization BUET
XNLI Conneau et al. (2018) 2018 15 NLI NYU
XQuAD Artetxe et al. (2020) 2020 10 Question Answering University of the Basque Country
XSEMPLR Zhang et al. (2023f) 2023 22 $\times$ Semantic Parsing PSU
XStoryCloze Lin et al. (2022) 2022 11 Commonsense Reasoning Meta
XTREME Hu et al. (2020) 2020 40 Mixed CMU
XTREME-R Ruder et al. (2021b) 2021 50 Mixed DeepMind
XWinograd Tikhonov and Ryabinin (2021) 2021 6 $\times$ Commonsense Reasoning Yandex
XCSR Lin et al. (2021c) 2021 16 Commonsense Reasoning USC
@@
Table 11: An overview of multilingual benchmarks with more than four supported languages after 2018 in which mBERT Devlin et al. (2018) was proposed. NLU and NLI denote the natural language understanding and the natural language inference task, respectively.
@@
With the emergence of novel models and algorithms, researchers inevitably scrutinize the capabilities by evaluating their performance on specific and challenging tasks Chang et al. (2024). Recently, LLMs have garnered significant interest in the academia and industry. The remarkable performance of LLMs has been proven on a variety of tasks, showing strong universality compared with prior PLMs limited to specific tasks. As shown in Table 11, to the best of our knowledge, we list the representative multilingual benchmarks after 2018 in which mBERT Devlin et al. (2018) was proposed. From the statistical results, we can observe many kinds of existing benchmarks, but there remain some issues with these benchmarks described as follows.

### v2

@@ 7.2.1 Legal LLMs in Multilingual Scenarios
Similar to the medical domain, the applications of LLMs in the legal domain principally center on English. Several precedent attempts such as Chatlaw [436], Lawyer LLaMA [448], SaulLM [449], and LegalBERT [450] expand the general ability of LLMs to the legal domain. The universal performance degradation has been observed when expanding to other languages.
To address the specific problems in the legal domain, the proposed models need to adapt the legal features, which are factual, ambiguous, structured, and timely [451, 452], compared to other domains. Brugger et al. [162] take a preliminary step in the multilingual legal Sentence Boundary Detection (SBD) covering 6 languages. Christen et al. [163] conduct similar research on the multilingual Negative Scope Resolution (detecting words affected by negation cue) task and Baumgartner et al. [164] extend legal judgment prediction to German, French and Italian. Furthermore, for a comprehensive evaluation, Niklaus et al. [161] collect 11 natural language understanding legal datasets covering a total of 24 languages and 8 subdivisions, which is the first cross-lingual legal benchmark (LEXTREME) showing there is still improvement room even for the popular models like ChatGPT. After the ChatGPT series models were proposed, Nguyen et al. [453] implement a preliminary empirical comparison between ChatGPT and GPT-4 on the “COLIEE” benchmark [454, 455] encompassing Japanese and English, where GPT-4 consistently surpasses its predecessor but still falls behind human performance.
For a more in-depth exploration of LLMs in multilingual legal issues, Niklaus et al. [165] construct a multilingual legal-domain corpus with 689GB, MultiLegalPile, whose detail is shown in Table 9. They train two PLMs based on XLM-R [447] and the Longformer [322] with this corpus, achieving the state-of-the-art performance on the LEXTREME [161] compared to XLM-R in most languages. Moreover, the English version of the model achieves state-of-the-art performance on 5 out of 7 tasks in LexGLUE [456], underscoring the exceptional cross-lingual ability and legal-domain expertise. Trautmann et al. [167] focus on employing legal prompt engineering (LPE) to enhance the capabilities of LLMs, thereby mitigating the challenges posed by the scarcity of cross-lingual legal data and the substantial computational resources required. Although improvements over baselines are observed, bridging the performance gap to match the supervised methods remains a significant endeavor.
@@ 7.3 Limitations and Future Directions on Multi-Domain
Data scarcity and translation issues. A powerful multilingual LLM, especially in specific domains, is predominantly hindered by the scarcity of domain data. Although knowledge transfer provides some relief, the issue of under-representation persists, particularly for low-resource languages, and becomes further compounded when extending to specific domains. Machine translation techniques offer a potential solution to mitigate this issue, however, they lead to new challenges [457]. On one hand, machine translation systems introduce errors, particularly when handling domain-specific terminology across multiple languages. Terms or phrases that native speakers do not use may also be included in the translated corpus. On the other hand, the models suffer from comprehensively understanding and accounting for the local and cultural context of the target language, complicating the task of in-depth and high-level feature learning and capturing.
@@ 8 Multilingual Data Resource
Amazon intent [168] 2022 49 2.02 GB Amazon Amazon
Amazon reviews [169] 2020 6 - Amazon Amazon
Aya [170] 2024 114 156 GB Mixed Cohere
Bactrian-x [171] 2023 51 - - MBZUAI
Biblenlp [172] 2024 861 581 MB Bible -
Bloom-lm [18] 2022 364 - - SIL International
CC100 [173] [174] 2020 109 185 GB CC Facebook
CulturaX [175] 2023 167 27 TB mC4, OSCAR University of Oregon
GPT-4 Prompts [176] 2024 5 1.05 GB GPT-4 -
Guanaco [177] 2023 4 400 MB - University of Washington
HPLT [178] 2023 75 22 TB CC, Internet Archive HPLT project
IWSLT 2017 [179] 2017 10 4.24 GB TED Talks FBK
mC4 [180] 2019 101 9.7 TB CC Google
Mewsli-x [181] 2021 50 285 MB WikiNews and Wikipedia DeepMind
Minds14 [182] 2021 12 471 MB Banking Assistant PolyAI Limited
MLDR [184] 2024 13 - Wikipeida, Wudao, mC4 BAAI
MMedC [185] 2024 6 105 GB Multiple SJTU
MQA [186] 2021 36 122 GB CC CLiPS Research
Multi-sentiments [187] 2022 12 141 MB Multiple Sources -
Multiconer [188] [189] 2023 12 338 MB Wikipedia Amazon
Open Subtitles [191] 2023 58 273 MB Subtitles -
OSCAR [192] 2020 166 6.3 TB CC University of Orego
Para-pat [193] 2020 15 2.57 GB Google Cloud University of Sheffield
Project Gutenberg [194] 2023 11 14.4 GB eBook Project Gutenberg
ShareGPT [195] 2023 5 1.08 GB GPT / Human Ronso
SREDFM [196] 2023 18 8.29 GB Wikidata, Wikipedia Babelscape
TED Talks [197] 2018 55 - TED Talks CMU
TED-talks-iwslt [198] 2012 104 25 GB TED Talks FBK
Toxi-text [199] 2023 55 1.96 GB - -
UD [200] 2023 102 2.19 GB - Universal Dependencies
Wikiann [201] [202] 2019 173 143 MB Wikipedia RPI
Wikipedia [203] 2024 322 71.8 GB Wikipedia Wikimedia Foundation
Wit Base [204] 2021 105 5.15 GB Wikipedia Google
xP3 [205] 2022 277 - - Cohere For AI
@@
LLMs are data-driven, thus the impressive learning capabilities of LLMs stem from their massive model sizes and extensive training datasets, which have been proven in high-resource languages. However, English stands as the closest approximation to the lingua franca, wielding dominance across various domains. With the largest number of total speakers, its prominence extends far and wide [458], where English reigns as the primary language of the internet [459, 460]. Meanwhile, English is the main language used in the higher economic status countries of the world, such as America, Britain, and other Western countries [461]. Therefore, existing data resources focus on the English-centric, which came at the expense of regional and indigenous languages, exacerbating language endangerment and economic marginalization [462]. Due to the lack of resources, this situation deeply restricts the development of multilingual models, and it is a vicious circle [463]. Moreover, low-resource languages suffer from lower quality, due to mislabeling or inadequate representation of native usage. This situation is especially prevalent with web-crawled data which predominantly consists of pornographic, nonsensical, or non-linguistic content [464].
@@
As shown in Table 10, we collect as much large-scale multilingual data resource as is reliable. We can observe that the scale of existing multilingual resources is much smaller than that of English monolingual resources (only four datasets with the TB level). And the low-/medium-resource languages typically derive data from a narrower range of sources compared to their high-resource counterparts [465]. The available data mainly originates from sources, such as Wikipedia, the Bible, and Common Crawl. In addition, these data suffer from bias and fairness, which we will present in Section 10. For instance, a gender bias exists in Wikipedia, with studies revealing a persistently low percentage of women editing articles [466]. The vicious circle of multilingual data includes the issue of open source issues. Due to the high cost of high-quality multilingual resources, researchers and companies are reluctant to share resources in the open-resource community (i.e., “data island”) [467, 468, 469, 470]. This situation results in multilingual research staying at the data level and ignoring the competition in the model paradigm. The united governments, companies, and researchers must start a virtuous cycle of multilingual data resources. Access to abundant, meticulously collected datasets in a language empowers researchers and developers to construct models and benchmarks. The abundance of models and benchmarks, in turn, fosters increased publication, facilitates communication, and promotes real-world application scenarios for companies. These outputs have the potential to attract more users, while government-mandated guidelines help generate non-toxic data, which can be repurposed for further research and development.
@@ 9 Multilingual Benchmark and Evaluation
M-Hellaswag [220] 2023 35 Commonsense NLI University of Oregon
XNLI [256] 2018 15 NLI NYU
Multilingual-Fig-QA [235] 2023 7 $\times$ NLI CMU
NoMIRACL [190] [82] 2023 16 $\times$ RAG University of Waterloo
MIRAGE-Bench [471] 2024 19 $\times$ RAG University of Waterloo
Cross-Sum [211] 2021 45 Summarization BUET
XL-SUM [255] 2021 44 $\times$ Summarization BUET
Pmindiasum [240] 2023 14 Summarization IIIT Hyderabad
SEAHORSE [242] 2023 6 $\times$ Summarization DeepMind
M3LS [222] 2023 20 $\times$ Summarization IITs
BELEBELE [208] 2023 122 Question Answering Meta
BioInstructQA [209] 2024 7 Question Answering Avignon Université
MLQA [231] 2020 7 Question Answering Facebook
TyDiQA [248] 2020 11 $\times$ Question Answering Google
XOR-TyDi [386] 2021 11 $\times$ Question Answering University of Washington
XQuAD [257] 2020 10 Question Answering University of the Basque Country
MaXM [227] 2023 7 Visual Question Answering Google
CIRAL [393] 2024 4 $\times$ Information Retrieval University of Waterloo
MIRACL [183] 2023 18 - Information Retrieval University of Waterloo
LAReQA [472] 2020 11 Information Retrieval Google Research
STSB-multi-mt [245] 2021 10 Text Similarity -
MasakhaNER [224] 2021 10 Named Entity Recognition Masakhane
Multi-CoNER [233] 2022 11 $\times$ Named Entity Recognition Amazon
XCOPA [251] 2020 12 Commonsense Reasoning Cambridge University
XCSQA [252] 2021 11 Commonsense Reasoning USC
XStoryCloze [38] 2022 11 Commonsense Reasoning Meta
XWinograd [261] 2021 6 $\times$ Commonsense Reasoning Yandex
XCSR [262] 2021 16 Commonsense Reasoning USC
FLORES-200 [215] [216] [217] 2022 200 Machine Translation Meta
OPUS-100 [238] 2020 100 Machine Translation University of Edinburgh
Tatoeba-mt [246] 2020 93 Machine Translation Tatoeba.org
Humaneval-XL [219] 2024 23 Code Generation University of Copenhagen
ODEX [237] 2022 14 Code Generation CMU
MARC [223] 2020 6 $\times$ Text Classification Amazon
Masakhanews [225] 2023 16 $\times$ News Topic Classification Masakhane
MULTIEURLEXDOC [234] 2021 23 Legal Topic Classification University of Copenhagen
Sib200 [243] 2024 205 Topic Classification UCL
Afrisent [206] 2023 14 $\times$ Sentiment Analysis U.Porto
ASPEN [207] 2022 31 Story Planning Cambridge University
Crossmodal-3600 [212] 2022 36 Image Captioning Google
Exams [213] 2020 15 Examination Sofia University
Fairlex [214] 2022 5 $\times$ Fairness University of Copenhagen
GEOMLAMA [218] 2022 5 Knowledge Diversity University of California
M-MMLU [220] 2023 35 NLU University of Oregon
M3Exam [221] 2023 9 $\times$ Examination DAMO Academy
MASSIVE [226] 2022 51 Intent Recognition, Slot Filling Amazon
Mela [230] 2022 10 $\times$ Linguistic Acceptability SJTU
MGSM [66] 2022 11 Mathematical Reasoning Google
MMedBench [232] 2023 6 $\times$ Medical SJTU
Paws-X [239] 2019 7 Paraphrase Identification Google
SMiLER [244] 2021 14 $\times$ Entity and Relation Extraction Samsung
Tydip [247] 2022 9 $\times$ Identify Politeness Levels UT-Austin
X-CLAIM [249] 2023 6 $\times$ Realworld Claims MBZUAI
X-RiSAWOZ [250] 2023 6 Dialogue Utterances Stanford University
xDial-Eval [253] 2023 10 Dialogues NUS
PRESTO [241] 2023 6 $\times$ Conversational Parsing University of Rochester
Universal Dependencies [200] 2022 146 $\times$ Parser -
XSEMPLR [258] 2023 22 $\times$ Semantic Parsing PSU
Bucc-bitext-mining [210] 2022 5 $\times$ Mixed HuggingFace
MEGA [228] 2023 70 $\times$ Mixed UW
MEGAVerse [229] 2023 83 $\times$ Mixed Microsoft
NusaX [236] 2023 10 Mixed Bloomberg
XGLUE [254] 2020 - $\times$ Mixed Microsoft
XTREME [259] 2020 40 Mixed CMU
XTREME-R [260] 2021 50 Mixed DeepMind
@@
Table 11: An overview of multilingual benchmarks with more than four supported languages after 2018 in which mBERT [271] was proposed. NLU and NLI denote the natural language understanding and the natural language inference task, respectively.
@@
With the emergence of novel models and algorithms, researchers inevitably scrutinize the capabilities by evaluating their performance on specific and challenging tasks [473]. Recently, LLMs have garnered significant interest in the academia and industry. The remarkable performance of LLMs has been proven on a variety of tasks, showing strong universality compared with prior PLMs limited to specific tasks. As shown in Table 11, to the best of our knowledge, we list the representative multilingual benchmarks after 2018 in which mBERT [271] was proposed. From the statistical results, we can observe many kinds of existing benchmarks, but there remain some issues with these benchmarks described as follows.